kylinn:

A computer downloaded 20 kilobytes of data in 4 seconds. If it downloads data at a constant speed, it can download kilobytes of data in 15 seconds.

4evablood:

whats the question

kylinn:

that is the question

4evablood:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kylinn that is the question \(\color{#0cbb34}{\text{End of Quote}}\) ok gimme a sec

4evablood:

75

Convert:

20/4= 5 kilobytes per second 15*5 = 75 kilobytes in 15 seconds

4evablood:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert 20/4= 5 kilobytes per second 15*5 = 75 kilobytes in 15 seconds \(\color{#0cbb34}{\text{End of Quote}}\) i said tht already

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @4evablood \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert 20/4= 5 kilobytes per second 15*5 = 75 kilobytes in 15 seconds \(\color{#0cbb34}{\text{End of Quote}}\) i said tht already \(\color{#0cbb34}{\text{End of Quote}}\) You can't give direct answers.

4evablood:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @4evablood \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert 20/4= 5 kilobytes per second 15*5 = 75 kilobytes in 15 seconds \(\color{#0cbb34}{\text{End of Quote}}\) i said tht already \(\color{#0cbb34}{\text{End of Quote}}\) You can't give direct answers. \(\color{#0cbb34}{\text{End of Quote}}\) sorry i didnt know

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @4evablood \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @4evablood \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert 20/4= 5 kilobytes per second 15*5 = 75 kilobytes in 15 seconds \(\color{#0cbb34}{\text{End of Quote}}\) i said tht already \(\color{#0cbb34}{\text{End of Quote}}\) You can't give direct answers. \(\color{#0cbb34}{\text{End of Quote}}\) sorry i didnt know \(\color{#0cbb34}{\text{End of Quote}}\) No worries! :)

4evablood:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @4evablood \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @4evablood \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert 20/4= 5 kilobytes per second 15*5 = 75 kilobytes in 15 seconds \(\color{#0cbb34}{\text{End of Quote}}\) i said tht already \(\color{#0cbb34}{\text{End of Quote}}\) You can't give direct answers. \(\color{#0cbb34}{\text{End of Quote}}\) sorry i didnt know \(\color{#0cbb34}{\text{End of Quote}}\) No worries! :) \(\color{#0cbb34}{\text{End of Quote}}\) ok